Consider a proper fraction such that the numerato is 2 less than the denominator. If the numerator and denominator are each increased by 1, then the ration of the numerator and denominator of the new fraction is 4:5. Determine the original fraction.
Answer by Theo(13342)

the original fraction is (x-2) / x
increase the numerator and denominator by 1 and you get (x-1) / (x+1)

you get (x-1) / (x+1) = 4/5
cross multiply to get 5*(x-1) = 4*(x+1)
simplify to get 5x-5 = 4x+4
subtract 4x from both sides of the equation and add 5 to both sides of the equation to get 5x-4x = 4+5
simplify to get x = 9

that's your answer.
the original fraction is 7/9.
add 1 to the numerator and denominator to get 8/10.
that simplifies to 4/5.

your solution is that the original fraction is 7/9.